Circular No. 8558

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 9P/TEMPEL C. W. Hergenrother, J. R. Weirich, and J. Keller, Lunar and Plane-

tary Laboratory, University of Arizona, report the results of aperture photometry on the inner coma of comet 9P/Tempel with the Kuiper 1.54-m reflector via a series of 30- and 60-s R-band integrations. The inner coma experienced a rapid brightening immediately after the impact by the ‘Deep Impact’ spacecraft onto the comet’s surface (around July 4.244 UT). The brightening began to level off ≈ 15 min after impact. By the time the comet was no longer observable, 70 min after impact, the brightening appeared to have plateaued. The following brightness increases were observed in apertures of varying diameters: 2′′.6, 2.15 mag; 3′′.5, 2.04 mag; 4′′.4, 1.92 mag; 5′′.2, 1.82 mag; 6′′.1, 1.74 mag; 7′′.0, 1.67 mag. The brightening was confined to the seeing disk of the pseudo-nucleus.

J. McGaha, Tucson, AZ, reports that comet 9P was observed with a low-light PC-164 video camera (exposing at 30 frames/s; spectral sensitivity peaking around 500–600 nm) attached to his 0.20-m f/4 reflector during July 4.240-4.257 UT; no flash was observed around the time of impact by the ‘Deep Impact’ spacecraft. Also, 90-s CCD images taken with a 0.62m reflector during July 4.240–4.302 show no changes, the coma and tail showing very well in these images.

Total visual magnitude estimates by M. Linnolt (Woodside, CA, 0.20m reflector), while the coma diameter remained steady at 2′.2: July 4.2368 UT, 11.2; 4.2451, 11.2; 4.2500, 11.1; 4.2583, 10.9; 4.2674, 10.7.

M. Kidger, Instituto de Astrofísica de Canarias, reports that his “Observadores cometas” group has obtained almost 2000 CCD photometric points of the light curve of 9P/Tempel since October 2004. There is a strong suggestion of a periodic rapid rise and slow decline, with the bestfit period being ≈ 4.40 ± 0.05 days, in both the light curve and in dust production measured as Afρ, that is presumably precessional.

COMET P/2005 M1 (CHRISTENSEN) Additional astrometry and the following improved elliptical orbital el-

ements appear on MPEC 2005-N20.

T = 2006 Jan. 31.809 TT  ω = 226.^{o}708
e = 0.30668  Ω = 143.046 }2000.0
q = 2.91826 AU  i = 10.149
a = 4.20910 AU n^{o} = 0.114135 P = 8.64 years
2005 July 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green